OpenStudy (anonymous):

Calculate the probability of drawing two cards from a 52-card deck. What is the probability of drawing a 2, replacing the card, and then drawing a King? a. 0.0044 b. 0.0045 c. 0.0060 d. 0.0059

OpenStudy (anonymous):

how many 2's in the deck?

OpenStudy (anonymous):

It does not say!

OpenStudy (anonymous):

i guess i should rephrase the question do you know how many 2's there are in a standard deck of cards?

OpenStudy (anonymous):

if you do not know this, you cannot answer the question better google standard deck of cards so fine out how many twos there are, and also how many cards total

OpenStudy (anonymous):

26?!

OpenStudy (anonymous):

no, only 4 two one of each suit \(\heartsuit, \diamondsuit, \spadesuit,\clubsuit\) hearts, diamonds, spades, clubs

OpenStudy (anonymous):

there are 4 twos, and 52 cards all together, so the probability you draw a 2 on the first pick is \(\frac{4}{52}=\frac{1}{13}\)

OpenStudy (anonymous):

there are also four kinds one of each suit the probability you pick a king is also \(\frac{4}{52}=\frac{1}{13}\) since you want the probability that both things happen, you take those two numbers and multiply them i.e. compute \[\frac{1}{13}\times \frac{1}{13}\]

OpenStudy (anonymous):

do I have to do cross multiply? or do i have to multiply straight across?